OpenStudy (anonymous):

Find n such that the error in the approximation of the definite integral is less than .00001 using the trapezoidal rule.

OpenStudy (anonymous):

∫011÷(1+x)dx

OpenStudy (anonymous):

sorry! thats the integral from 0 to 1 of 1/(1+x)dx

OpenStudy (anonymous):

I just don't know how to use the max / abs value part of the formula

OpenStudy (anonymous):

the integral from 0 to 1 of 1/(1+x)dx f(x)=1/(1+x), f '(x)=1/(1+x)^2, and f ''(x)=2/(1+x)^2, f ''(0)=2/(1+0)^2=2 now 10^-5 <[(1-0)/12]h^2 M, M=2, but h=1/n, [1/12](1/n)^2 M <10^-5 1/(6n^2)<10^-5 10^5 /6< n^2 sqrt(10^5 /6)< /n/ 129.1<n use n=130 and beyond answer